OpenStudy (anonymous):

1/4y+3=-5x / a formula and function problem

OpenStudy (anonymous):

This is posted again after...

OpenStudy (anonymous):

Rearrange it?

jhonyy9 (jhonyy9):

1+12y=-20xy x=(1+12y)/-20y= -(1+12y)/20y or for y 12y+20xy=-1 y(12+20x)=-1 y=-1/(12+20x)